OpenStudy (anonymous):

Factor. 8x^3+14x^2-15x

OpenStudy (campbell_st):

start by taking out an x as a common factor \[x(8x^2 + 14x - 15)\] next you need to multiply 8 and -15 = -120 find the factors that add to 14, the larger is positive so 20 and - 6 then you get \[\frac{x(8x + 20)(8x - 3)}{8}\] remove the common factors from the binomials \[\frac{x \times 4(2x + 5)\times 2(4x -3)}{8}\] cancel the common factors leaving x(2x + 5)(4x - 3)

OpenStudy (anonymous):

Thank You!